OpenStudy (anonymous):

When converting a measurement in cm to a measurement in km. For example 20.7 cm when I switch that to decimeters its 2.07 dm then .207 m. Then deka meter .0207 then hecto .00207 then kilometer .000207 km 2.07e-4 km However my book says its 2.07e-9 km. How is that happening? What am I missing?

OpenStudy (anonymous):

You are correct according to your post. You sure the problem is not asking you to convert \[20.7cm^2 \to km^2\]

OpenStudy (anonymous):

That is what its asking. lol I figured that the squared didn't matter. Seeing as both are squared I thought I could deal with the numbers independently.

OpenStudy (anonymous):

So im guessing that the squared has something to do with why im confused but what about it being squared changes my answer?

OpenStudy (anonymous):

cm2=(cm)(cm) and km2=(km)(km) Start with given variable and units Solve for the unknown units. Try to make the starting units cancel out by multiplying the cm in denominator twice. You want km^2 so multiply the km twice by putting it as a numerator over the cm. So it equals to this \[\frac{ 20.7cm^2 }{ 1 }\times \frac{ ?km }{ 1cm }\times \frac{ ?km }{ 1cm }\] Fraction and numerator should equal after starting out the given variable. Since 0.00001km=1cm then it \[\frac{ 20.7cm^2 }{ 1 }\times \frac{ 0.0001km }{ cm }\times \frac{0.0001km }{ cm }\] Multiply it and the cm units would cancel out leaving you with just the km^2 and it would be 0.000000000207km or 2.07 x 10^-9km or 2.07e^-9

OpenStudy (anonymous):

20.7 * 0.0001 * 0.0001 = 0.000000207 = 2.07e-7 How am I still missing this?

OpenStudy (anonymous):

I made a typo, supposed to be 20.7km*0.00001km*0.00001km= I made no typo explaining,just in the conversion example.